Hints ion w= 13.0 J of Submit My Answers Give Up IS Correct The energy transferred to the tomato by doing work on it increases its potential energy because the tomato is now at a greater height or position relative to the ground. used Kinetic energy Kinetic energy, Ek is gr en by he formula醜 1 mu2 where m is mass s he mass in lo rams and u e ct n meter per e con Part B The tomato s dropped. What is the velocity. U of the tomato when it hits the ground? Assume 947 % of the work done in Part A s transferred to kinetic energy, E, by the time the tomato hits the ground Express your answer with the appropriate units Hints v= 11.51

The weight of the tomato is 105 g.
Please work problem. I can't figure out what I'm doing wrong
Answer #1

Given , m = 105 g = 0.105 kg , W = 13 J

Since 94.7% of work done is converted to Kinetic energy(Ek)

Ek = (94.7/100)*13 = 12.311 J

So,

Ek = (1/2)*m*v^2

12.311 = (1/2)*0.105*v^2

v^2 = 234.495

v = 15.313 m/s

Thus, velocity of tomato is 15.313 m/s